# Analysis of a voter model with an evolving number of opinion states

### v1

@@ I Introduction
The original VM has been extended in many different ways. This includes variants of the VM with ‘noise’, allowing individuals to change opinion states spontaneously [25, 26, 27], nonlinear voter models [28, 29, 30], multistate voter models (models with more than two opinion states) [31, 32, 33], the coevolving voter model [34, 35, 36] in which agents can re-wire links in the interaction network, and higher-order voter models [37].
@@ II.3 Absence of absorbing states
Figure 2: (a) The distribution $P(M)$ of the number of opinions in the steady state is shown for different values of $\alpha$. Numerical results (symbols) for $N=2^{12}$, averaged over $5000$ samples and the theoretical predictions (lines) from Eq. (2) are shown together. (b) Markers show the average number of opinions, $M$ as a function of $\alpha N$ for various population sizes $N$ as indicated. Lines show again the theoretical predictions from Eq. (2). The inset shows a magnification for small $\alpha$ with various $N$ with the same symbols. The vertical lines in the inset represent the location of $\alpha N=1/\log N$.
@@ III.1.2 Distribution of the number of opinions
Using an analogy to models of the evolution of mating types in [32] we can calculate the distribution $P(M)$ of the number of different opinions present in the population in the stationary state. More specifically, the imitation mechanism in the VM is akin to asexual reproduction in the evolutionary model in [32]. In an event of this type an individual reproduces and generates an offspring which is of the same type as the parent. To keep the population size constant, a random other individual is removed from the population following a reproduction. Such a combined reproduction-removal event has the same effect as an imitation event in our model. Similarly, the setup in [32] allows for mutation, which is modelled as spontaneous state change of an individual to a mating type that is assumed not to be present in the population. This is analogous to the introduction of new opinions in the VM. For further details see in particular Fig. 1 in [32], and the text in Sec. 2.1 of that paper.
@@
where $\begin{bmatrix}N\\ M\end{bmatrix}$ is the unsigned Stirling number of the first kind. This prediction is successfully tested against simulations for various values of $\alpha$ in Fig. 2(a). From this distribution we can directly calculate the mean number of opinions in the population in the stationary state.
@@ III.1.3 Typical number of opinions in the population
The average number of opinions $M$ in the steady state is shown in Fig. 2(b) as a function of $\alpha N$ for different population sizes. The number of distinct opinions grows with the rate $\alpha N$. The model exhibits two distinct behaviors: for small values of $\alpha$, the system is near consensus ($M\approx 1$). In this regime ($\alpha N\ll 1$), the introduction of new opinions is rare and if any new opinion emerges, it typically goes extinct before a further new opinion can appear. Conversely, for $\alpha N\gg 1$, the number $M$ of opinions increases with increasing $\alpha N$.
@@ III.2.1 Distribution of number of agents holding a particular opinion
What we really measured in the simulations described in first paragraph of this Sec. III.2.1 is the distribution of $s$ conditioned on the fact that the opinion is present in the system (i.e., that $s>0$). This object is, for $s>0$,
@@ III.3 Fraction of time spent in the consensus state
Following on from Secs. III.1.3 and III.2 we study the fraction of time the system remains in consensus as a function of $\alpha N$. Specifically, we measured the total time spent in any of the consensus states, relative to the total simulation time (in the stationary state), i.e., $t_{\rm cons}/t_{\rm tot}$. This quantity is the probability to find the system in consensus at any random time in the stationary state, given by $P(M=1)$ in Eq. (2). By examining this fraction for different values of $\alpha N$, we can identify the conditions under when a consensus becomes possible with non-vanishing probability, and when it becomes the prevailing outcome.
@@ III.4 Density of active interfaces
We next calculate the time evolution and asymptotic value of the density of active links in the system. Our calculation assumes an infinite system size, but other than that no approximation is made. The density of active links is the proportion of links connecting agents who are in different opinion states. We write $\rho(t)$ for this density at time $t$.
@@
Using the relations
@@
with $\rho^{*}=\frac{\alpha(N-1)}{1+\alpha(N-1)}$ and $\delta=\rho^{*}/(2\alpha)$. Thus, $\rho$ decays towards the asymptotic value $\rho^{*}$, with decay time $\delta$. The density $\rho^{*}$ of active links in the steady state is non-zero unless $\alpha=0$ (in this latter case the model becomes the identical to the conventional VM). We note that $\rho^{*}=1/2$ when $\alpha(N-1)=1$; for large $N$ this is the same as the condition $\alpha N=1$.

### v2

@@ I Introduction
The original VM has been extended in many different ways. This includes variants of the VM with ‘noise’, allowing individuals to change opinion states spontaneously [25, 26, 27], nonlinear voter models [28, 29, 30], multistate voter models (models with more than two opinion states) [31, 32, 33, 34, 35, 36, 37], the coevolving voter model [38, 39, 40] in which agents can re-wire links in the interaction network, and higher-order voter models [41].
@@ II.3 Absence of absorbing states
Figure 2: (a) The distribution $P(M)$ of the number of opinions in the steady state is shown for different values of $\alpha$. Numerical results (symbols) for $N=2^{12}$, averaged over $5000$ samples and the theoretical predictions (lines) from Eq. (2) are shown together. (b) Markers show the average number of opinions, $M$ as a function of $\alpha N$ for various population sizes $N$ as indicated. Lines show the theoretical predictions obtained by numerically evaluating the first moment of the distribution in Eq. (2). The inset shows a magnification for small $\alpha$ with various $N$ with the same symbols. The vertical lines in the inset represent the location of $\alpha N=1/\log N$. The dashed line in panel (b) indicates power-law increase with an exponent of $0.85$ (obtained from a fit).
@@ III.1.2 Distribution of the number of opinions
Using an analogy to models of the evolution of mating types in [35] we can calculate the distribution $P(M)$ of the number of different opinions present in the population in the stationary state. The imitation mechanism in the VM is akin to asexual reproduction in the evolutionary model in [35]. In an asexual reproduction event an individual reproduces and generates an offspring which is of the same type as the parent. To keep the population size constant, a random other individual is removed from the population in [35] following a reproduction. If this removed individual is of a type different to that of the reproducing individual such a combined reproduction-removal event has the same effect as an imitation event in our model – an agent of one type (opinion) is replaced by an agent of another type. The rate for such events is indicated in Table I in [35], and is in direct analogy to step 2 of the update algorithm in Sec. II.1 of the current paper. Similarly, the setup in [35] allows for mutation, which is modelled as spontaneous state change of an individual to a mating type that is assumed not to be present in the population. This is analogous to spontaneous change of an agent to a new opinion not currently represented by any other agent. For further details see in particular Fig. 1 in [35], and the text in Sec. 2.1 of that paper.
@@
where $\begin{bmatrix}N\\ M\end{bmatrix}$ is the unsigned Stirling number of the first kind. This prediction is successfully tested against simulations for various values of $\alpha$ in Fig. 2(a). From this distribution we can directly calculate the mean number of opinions in the population in the stationary state. Although the binomial coefficient in Eq. (2) is formally defined only for integer arguments, we extend the equation to non-integer values of $N\alpha$ via the gamma function representation of factorials. In this way our result can be evaluated for non-integer values of $N\alpha$
@@ III.1.3 Typical number of opinions in the population
The average number of opinions $M$ in the steady state is shown in Fig. 2(b) as a function of $\alpha N$ for different population sizes. The number of distinct opinions grows with the rate $\alpha N$. The model exhibits two distinct behaviors: for small values of $\alpha$, the system is near consensus ($M\approx 1$). In this regime ($\alpha N\ll 1$), the introduction of new opinions is rare and if any new opinion emerges, it typically goes extinct before a further new opinion can appear. Conversely, for $\alpha N\gg 1$, the number $M$ of opinions increases approximately as a power law with increasing $\alpha N$.
@@ III.2.1 Distribution of number of agents holding a particular opinion
This expression applies for $s=0,1,\dots,N$ (for $s=0$ the numerator takes the value of one). The denominator ensures normalisation. What we really measured in the simulations described in first paragraph of this Sec. III.2.1 is the distribution of $s$ conditioned on the fact that the opinion is present in the system (i.e., that $s>0$). This object is, for $s>0$,
@@ III.3 Fraction of time spent in the consensus state
Following on from Secs. III.1.3 and III.2 we study the fraction of time the system remains in consensus as a function of $\alpha N$. Specifically, we measured the total time spent in any of the consensus states, relative to the total simulation time (in the stationary state), i.e., $t_{\rm cons}/t_{\rm tot}$. This quantity is the probability to find the system in consensus at any random time in the stationary state, given by $P(M=1)$ in Eq. (2). Using the identity $\begin{bmatrix}N\\ 1\end{bmatrix}=(N-1)!$ we find
$P(M=1)=\frac{1}{\binom{N\alpha+N-1}{N\alpha}}.$ (9)
By examining this fraction for different values of $\alpha N$, we can identify the conditions under when a consensus becomes possible with non-vanishing probability, and when it becomes the prevailing outcome.
@@ III.4 Density of active interfaces
We next calculate the time evolution and asymptotic value of the density of active links in the system (we focus on all-to-all connectivity, thus the word link refers to connections between any two nodes). Our calculation assumes an infinite system size, but other than that no approximation is made. The density of active links is the proportion of links connecting agents who are in different opinion states. We write $\rho(t)$ for this density at time $t$.
@@
The notation $\sum_{a<b}\dots$ indicates a sum over $a$ and $b$, covering all pairs $a,b\in\{1,\dots,M\}$ such that $a<b$. Using the relations
@@
with $\rho^{*}=\frac{\alpha(N-1)}{1+\alpha(N-1)}$ and $\delta=\rho^{*}/(2\alpha)$. Thus, $\rho$ decays towards the asymptotic value $\rho^{*}$, with decay time $\delta$. The density $\rho^{*}$ of active links in the steady state is non-zero unless $\alpha=0$ (in this latter case the model becomes the identical to the conventional VM). From the expression for $\rho^{*}$ we note that $\rho^{*}=1/2$ when $\alpha(N-1)=1$. Fig. 6(b) indicates a cross-over of the stationary density of active interfaces from $\rho^{*}=0$ for small $\alpha N$ to $\rho^{*}=1$ for large $\alpha N$. Thus, $\rho^{*}$ takes the separating value $\rho^{*}=1/2$ when $\alpha(N-1)=1$. For large $N$ this reduces to $\alpha N=1$, and hence to the condition at which the quasi-stationary distribution $Q(s)$ changes shape [Eq. (7)]. The same condition also comes up in Sec. III.3.